OpenStudy (anonymous):

f(x) = x^2 - 16 and g(x) = x+4. Find f/g and its domain.

OpenStudy (anonymous):

Its domain is -4, but what is f/g?

OpenStudy (agent0smith):

Domain is all real numbers EXCEPT x=-4

OpenStudy (anonymous):

Yeah that ^ lol cx Idk how to get the rest its not the same as the last question cause it has an exponent

OpenStudy (anonymous):

x^2-16/x+4 how do i solve that ?

OpenStudy (agent0smith):

Use brackets, just like last time.

OpenStudy (anonymous):

I dont get it ...howw? /.\

OpenStudy (anonymous):

wait i tell you how to solve it..@SkiTTleoooo47

OpenStudy (anonymous):

@agent0smith is already helping me :)

OpenStudy (agent0smith):

lol... put brackets in :P like this: (x^2-16)/(x+4) Now you have to factor the x^2-16. Remember difference of two squares?

OpenStudy (anonymous):

The difference of two squares?

OpenStudy (anonymous):

\[f/g=\frac{ x^{2}-16 }{ x+4 } =\frac{ (x+4)(x-4) }{ (x+4) } =(x-4)\] in 2nd step use this formula...a^2-b^2=(a+b)(a-b)

OpenStudy (agent0smith):

difference of two squares:\[\large a^2 - b^2 = (a+b)(a-b)\]

OpenStudy (anonymous):

Oh okay okay lol Let me find another pen lol i messed up the one i had :P

OpenStudy (anonymous):

Wait do i put in for the value's of A and b?

OpenStudy (agent0smith):

well, \[\large x^2 - 16 = x^2 - 4^2\] right?

OpenStudy (anonymous):

Yeah ?

OpenStudy (agent0smith):

so then \[\large x^2 - 4^2 = (x+4)(x-4)\]since\[\large a^2 - b^2 = (a+b)(a-b)\]

OpenStudy (anonymous):

So what do i do next i solve that?

OpenStudy (agent0smith):

Now that we factored it, \[\Large f/g=\frac{ x^{2}-16 }{ x+4 } =\frac{ (x+4)(x-4) }{ (x+4) } \]now try simplifying

OpenStudy (anonymous):

f/g= x^2/-16/x+4= (x+4) ?

OpenStudy (agent0smith):

Look again.

OpenStudy (anonymous):

Where am i wrong?

OpenStudy (agent0smith):

look again at my post, and you might see ;P

OpenStudy (anonymous):

I dont see it! Omg :( Why was i born so dumbbb lol

OpenStudy (agent0smith):

\[\Large \frac{ (x+4)(x-4) }{ (x+4) }\]simplify this.

OpenStudy (agent0smith):

What cancels off on the top and bottom?

OpenStudy (anonymous):

x-4 and x+4 right? or is it x+4 and x+4?

OpenStudy (agent0smith):

Cancel off common factors, whatever is EXACTLY the same on the top and bottom. What's left?

OpenStudy (anonymous):

x-4?

OpenStudy (agent0smith):

Correct.

OpenStudy (anonymous):

Okay so its f/g= x^2-16/x+4+(x-4)? right ?

OpenStudy (agent0smith):

What? You literally just worked out the answer a moment ago...

OpenStudy (anonymous):

i meant an equal sign lol

OpenStudy (agent0smith):

Tell me what f/g is, your last post is not it :P

OpenStudy (anonymous):

Oh so the answer was x-4?

OpenStudy (agent0smith):

f/g = x-4, domain is earlier somewhere up top

OpenStudy (anonymous):

-4? :)

OpenStudy (anonymous):

So my answer is f/g= x-4 and domain is all real numbers except x=-4?

OpenStudy (agent0smith):

Yep.

OpenStudy (anonymous):

Okay thanks :)

OpenStudy (agent0smith):

no prob :)